OpenStudy (anonymous):

identify the range of the equation y = −x2 − 8x − 10.

OpenStudy (btaylor):

since the x^2 term has a negative coefficient, we know that the parabola opens downward. Now, we solve for the vertex. The x-value of the vertex is -b/2a, so 8/-2 = -4. Now plug -4 in for x, and you get the y value at the maximum/vertex. -(-4)^2 - 8(-4) - 10 = 6. So the vertex is at (-4 , 6) If you were to graph this parabola, it would look roughly:The range is all possible y values, which is here from negative infinity to 6. In interval notation:\[\left\{ x \in \mathbb{R} | x \in (-\infty,6]\right\} \]

OpenStudy (anonymous):

nvm i do thank you